the price of a dishwasher today is $500 and inflation is 10% per year. therefore, in 3 years the price will be $650. true or false
Answer #1

Present value = PV = $500

inflation rate = i = 10% per year

Time period = n = 3 years

Future value can be calculated using the formula:

FV = PV*(1+i)n

Future value of dishwasher = FV = 500*(1+10%)3 = 500*1.13 = 500*1.331 = $665.50

Price of the dishwasher in 3 years will be $665.5 and not $650.

Answer -> the price of a dishwasher today is $500 and inflation is 10% per year. therefore, in 3 years the price will be $650 - False